OpenStudy (anonymous):

Find dy/dx by implicit differentiation: 4 cos x sin y = 1

OpenStudy (experimentx):

- 4 sin x sin y + 4 cos x cos y dy/dx = 0 => dy/dx = sin x sin y/ cos x cos y = tan x tan y

OpenStudy (lalaly):

use the product rule,, \[-4sinxsiny+4cosxcosy \frac{dy}{dx}=0\]\[\frac{dy}{dx}=\frac{4sinxsiny}{4cosxcosy}=tanxtany\]

sam (.sam.):

4cosxsiny=1 differentiate, 4[(cosx)(cosy)(y')+(siny)(-sinx)]=0 4(cosx)(cosy)(y')=4sinxsiny (y')=4sinxsiny /4(cosx)(cosy)

OpenStudy (anonymous):

thanks guyz